# Structural Cointegration of the Paleoclimate

### v2

@@ Abstract
This paper investigates the coupled dynamics of temperature and atmospheric CO2 concentrations during Pleistocene glacial cycles using a Structural Vector Error Correction Model (SVECM). To address the controversy regarding causal leads and lags, we develop a structural identification framework that utilizes Milankovitch orbital parameters as instrumental variables to disentangle the Instantaneous Greenhouse Effect (IGE) and the Instantaneous Carbon Feedback (ICF) at a millennial resolution. Focusing on the gap-free interval of Termination II, we employ a Causality Allocation Curve (CAC) derived from system-wide residuals to identify the causal structure that maximizes statistical consistency with the observed record. This approach allows for the estimation of Earth System Sensitivity (ESS) based purely on observational data and the decomposition of warming dynamics into structural momentum and equilibrium adjustment. By demonstrating the consistency of these paleo-derived structural parameters with modern instrumental records, this study establishes a robust econometric foundation for integrating long-term climate analysis with contemporary climate monitoring.
@@ Introduction
The strong coupling between atmospheric CO2 concentrations and global temperatures over the glacial-interglacial cycles of the Quaternary period is one of the most striking features of the Earth’s climate system. High-resolution ice-core records from Antarctica, spanning the last 800,000 years, have provided unequivocal evidence that CO2 and temperature covary with remarkable precision across multiple terminations and glaciations (Jouzel et al., 2007; Lüthi et al., 2008). This tight correlation forms the empirical basis for our understanding of the greenhouse effect as a fundamental regulator of the Earth’s long-term climatic state.
However, the precise causal architecture of this relationship remains a subject of intense debate in paleoclimatology. Early investigations into the timing of these transitions suggested that Antarctic temperature increases preceded CO2 rises by several centuries during glacial terminations, leading to the hypothesis that CO2 acts primarily as an amplifier of orbitally-induced warming rather than its primary driver (Caillon et al., 2003). Conversely, recent re-evaluations of ice-core gas-age chronologies have argued that the perceived lag may be a statistical artifact of the firn densification process, suggesting that CO2 and temperature are effectively synchronized at the millennial scale (Parrenin et al., 2013). This “lead-lag” controversy is not merely a chronological puzzle; it strikes at the heart of our ability to quantify the true climate sensitivity and the structural feedbacks inherent in the Earth system.
The fundamental difficulty in resolving this causality lies in the simultaneous nature of the climate-carbon coupling. Within the observational resolution of the paleoclimatic record, temperature and CO2 operate in a continuous feedback loop: warming triggers the release of CO2 from the oceans and terrestrial biosphere, while increased CO2 further warms the planet through radiative forcing. In econometric terms, this constitutes a problem of simultaneity and endogeneity, where standard correlation analyses or simple Granger-causality tests are insufficient to isolate the structural parameters of the system. Without a robust identification strategy, it is impossible to determine whether the observed covariance is driven by an instantaneous greenhouse effect or by a temperature-driven carbon response.
In this paper, we address this identification challenge by applying a Structural Vector Error Correction Model (SVECM) to the Pleistocene ice-core record. Our approach builds upon the methodology of structural macroeconometrics, utilizing the strict exogeneity of Earth’s orbital parameters—Milankovitch forcing—to achieve causal identification. We distinguish between two contemporaneous mechanisms operating within a 1,000-year time step: the Instantaneous Greenhouse Effect (IGE), representing the immediate thermal response to CO2 innovations, and the Instantaneous Carbon Feedback (ICF), representing the carbon cycle’s response to warming.
By analyzing the gap-free record of Termination II, we demonstrate that the Pleistocene climate is characterized by a null Instantaneous Greenhouse Effect ($a_{12}\approx 0$) at the millennial resolution. Instead, we show that the greenhouse effect is structurally localized as a powerful delayed feedback. Our model identifies a structural lag of 1,000 years, after which a CO2 doubling triggers a cumulative temperature increase of approximately 11.8 K, comprising a 10.5 K structural momentum and a 1.3 K long-run equilibrium adjustment.
Crucially, we validate these paleo-derived structural parameters by projecting them onto the modern era. We show that the centennial warming rate derived from our Pleistocene model accurately recovers the observed instrumental warming of the last 150 years. This consistency suggests that the fundamental coupling between the thermal and carbon cycles has remained stable across vastly different climatic regimes. By providing a definitive empirical verification of the structural lag in the greenhouse effect, this study offers a new framework for reconciling the millennial-scale dynamics observed in the ice cores with the centennial-scale warming observed in the modern anthropogenic era.
While earlier studies, such as Kaufmann et al. (2006), have successfully applied VECM to instrumental climate records to identify long-run relationships, they relied on pre-defined radiative forcing formulas and reduced-form estimations. By contrast, our study extends this econometric framework into the structural domain (SVECM). By utilizing the orbital forcing as a source of exogenous variation, we identify the distinct structural parameters of the Earth system—IGE, ICF, and the structural lag effects. This allows us to estimate an Earth System Sensitivity (ESS) that is derived purely from observational data, consistent with the long-term feedback framework proposed by the PALAEOSENS project (Rohling et al., 2012).
@@
To circumvent this risk while preserving the critical information contained in the levels of the variables, we employ a Vector Error Correction Model (VECM). This framework is appropriate when variables are cointegrated, meaning that while individual series may drift non-stationarily, a specific linear combination of them remains stationary. This stationary combination represents a long-run equilibrium or a physical bond that has persisted through numerous glacial-interglacial cycles. The VECM is formulated as follows:
@@
where $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$ is the vector of endogenous variables representing temperature ($T_{t}$), log-transformed CO2 concentration ($\ln C_{t}$), and orbital forcing ($M_{t}$). $\Delta$ denotes the first-difference operator, and $\epsilon_{t}$ is a vector of reduced-form residuals.
The matrix $\Pi=\alpha\beta^{\prime}$ contains the essential information regarding the cointegrating space. The vector $\beta$ defines the long-run equilibrium toward which the climate system gravitates over centennial to millennial scales. The existence of a stable $\beta$ suggests a fundamental consistency in the coupling of the thermal and carbon cycles across diverse climatic regimes.
The matrix $\alpha$ represents the speed of adjustment or the restoring force of the system. When a stochastic shock—such as an orbital perturbation—displaces the system from its equilibrium path, the error correction term ($\beta^{\prime}y_{t-1}$) becomes non-zero. The coefficient $\alpha$ dictates how rapidly the system initiates corrective dynamics. This mechanism constitutes the mathematical representation of the climate system’s self-regulation.
@@
To isolate structural causalities from the reduced-form correlations, we assume that the VECM is derived from an underlying Structural VAR (SVAR) model. This approach follows the framework established by Sims (1980), which provides a basis for analyzing endogeneity and structural innovations within a system of interrelated time series by utilizing the residuals of the estimated model. By applying this structural methodology to the paleoclimatic system, we can define a model that explicitly accounts for the contemporaneous interactions between temperature and carbon cycles:
$(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}1\Delta y{t-1}+\dots+u_{t},$ $u_{t}\sim\text{IID}(0,I)$ (2)
The matrix $A_{0}$ represents the simultaneous feedbacks within the same 1,000-year time step. The reduced-form residuals $\epsilon_{t}$ are related to the structural shocks $u_{t}$ via the relation:
$u_{t}=(I-A_{0})\epsilon_{t}$ (3)
Here, the elements $a_{ij}$ of the structural matrix $A=I-A_{0}$ designate the contemporaneous causal effects.
Following physical constraints, we assume that temperature and CO22 do not affect orbital forcing ($a{31}=a_{32}=0$), and that orbital forcing does not directly affect CO22 concentration (a23=0a{23}=0). Under these restrictions, Equation (3) yields the following system of structural innovations:
$\begin{aligned} u_{1t}&=\epsilon_{1t}-a_{12}\epsilon_{2t}-a_{13}\epsilon_{3t}\\ u_{2t}&=-a_{21}\epsilon_{1t}+\epsilon_{2t}\\ u_{3t}&=\epsilon_{3t}\end{aligned}$ (4)
where $a_{12}$ denotes the Instantaneous Greenhouse Effect (IGE)—the immediate temperature response to a change in CO22 within the same millennium—and a21a{21} denotes the Instantaneous Carbon Feedback (ICF). Importantly, $a_{12}=0$ corresponds to a null-IGE identification, where CO2 increases exert no additional warming effect at the contemporaneous resolution.
Identification of these parameters requires the orthogonality of structural shocks ($\mathbb{E}[u_{it}u_{jt}]=0$ for $i\neq j$). Substituting Equation (4) into the condition $\mathbb{E}[u_{1t}u_{2t}]=0$, we derive the Causality Allocation Curve (CAC):
@@
This hyperbola represents all causal allocations $(a_{12},a_{21})$ consistent with the observed residual covariances $s_{ij}$.
@@
To evaluate the dynamic stability and long-term consequences of structural shocks, we employ impulse response functions (IRFs). By utilizing the moving average representation of the VECM, the level of the variables at horizon $h$ is expressed as:
$y_{t+h}=\Psi_{0}u_{t+h}+\Psi_{1}u_{t+h-1}+\dots+\Psi_{h}u_{t}$ (6)
The matrix $\Psi_{j}$ represents how a structural innovation at time $t$ propagates through the simultaneous interactions and the long-run adjustment mechanism. We simulate two hallmark scenarios to quantify climate sensitivity:
@@
    A CO2 doubling shock: Represented as a structural innovation in the carbon cycle ($u_{2t}$). In our logarithmic specification, a doubling is equivalent to an instantaneous shock of magnitude $\ln(2)\approx 0.693$. This allows us to trace the temperature response via the IGE and subsequent feedback loops.
@@
    A 2K temperature shock: Represented as an abrupt warming ($u_{1t}=2.0$). This scenario quantifies the magnitude of the resulting CO2 release, identifying the strength of the ICF and the efficiency of the error correction process.
Our VECM is estimated on a 1,000-year grid to ensure a high signal-to-noise ratio in the ice-core proxies. While the structural dynamics are defined on this millennial scale, we utilize spline interpolation when presenting the IRFs to facilitate comparison with higher-frequency records without altering the fundamental parameters.
As derived in Appendix A, the presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline. The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$, represents the total accumulated effect of a shock. This property provides a bridge between millennial-scale dynamics and long-term climate transitions, ensuring that the projected trajectories remain robust across diverse time scales.
@@
Sources of Paleoclimate and Orbital Data
Figure 1: Evolution of climatic and orbital variables over the past 800,000 years. The panels display (from top to bottom) solar insolation at $60^{\circ}$N in June and in December, Antarctic temperature anomalies ($\Delta T$), and the logarithm of atmospheric CO2 concentration ($\ln C$), highlighting the strong phase-locking of climate variables to orbital cycles.
@@
The empirical foundation of this study rests on three primary data series spanning the late Quaternary: atmospheric CO2 concentrations, Antarctic surface temperature anomalies, and solar insolation. These datasets provide the necessary temporal coverage to examine the long-term cointegrating relationships and the millennial-scale structural dynamics of the Earth system.
Atmospheric CO2 data are obtained from the composite ice-core record provided by the World Data Service for Paleoclimatology (Lüthi et al., 2008; Bereiter et al., 2015). This high-resolution dataset integrates measurements from the EPICA Dome C, Vostok, and other Antarctic ice cores, extending back to 800,000 years before present (yr BP). This record captures the natural variability of greenhouse gases across eight complete glacial-interglacial cycles, providing a robust baseline for evaluating climate sensitivity.
For temperature estimates, we utilize the Antarctic temperature anomaly record derived from stable isotope (deuterium) ratios in the EPICA Dome C ice core (Jouzel et al., 2007). The deuterium record provides a consistent proxy for regional Antarctic temperature, which serves as a reliable indicator for global climatic shifts due to the strong coupling between polar and global temperature variations.
Solar insolation data, representing the external orbital forcing, are computed using the IMCCE online numerical integration service based on the solution by Laskar et al. (2004). Specifically, we utilize the mean daily insolation at $60^{\circ}$N during the summer solstice in June. This metric is a standard choice (adopted by Lemoine, 2010) for capturing the orbital drivers of glacial-interglacial transitions, commonly referred to as Milankovitch forcing. The raw series of these variables, which exhibit clear cyclicality and phase-locking, are plotted in Figure 1.
Data Gridding and the Selection of Termination II
The raw ice-core samples are characterized by uneven temporal spacing, which typically becomes sparser with increasing age. To conduct a consistent time-series analysis using the Vector Error Correction Model (VECM), the data must be transformed into a regular temporal grid. We adopt a two-stage gridding process designed to preserve the integrity of the original physical signal while ensuring statistical reliability.
In the first stage, the raw data are allocated into 100-year bins. When a bin contains multiple observations, the values are averaged to reduce sampling noise and provide a representative value for that centennial interval. In the second stage, these 100-year records are aggregated into a 1,000-year grid, which serves as our primary unit of observation ($t$). For each 1,000-year grid point, the representative value is calculated by averaging the available 100-year bin values within that specific millennium.
Crucially, for our primary structural analysis, we do not employ interpolation during this aggregation process. If a 1,000-year window contains no underlying 100-year data points, it is intentionally left as a gap. This approach ensures that the resulting 1,000-year series remains a direct reflection of the observed ice-core record, free from any statistical artifacts that could be introduced by infilling missing values. Linear interpolation at the 100-year level is utilized exclusively for the Whole Period analysis (spanning 800,000 years) to facilitate a continuous estimation over the entire record where data sparsity is more pronounced.
As illustrated in Figure 2, our grid analysis identifies the presence of observational gaps across the late Quaternary. The longest continuous span that is entirely free of such gaps at the 1,000-year resolution—and therefore requires no interpolation whatsoever—perfectly coincides with the period known as Termination II (approximately 141,000 to 121,000 yr BP). This period captures the rapid transition from the penultimate glacial maximum to the Eemian interglacial.
By focusing our primary structural analysis on this gap-free Termination II sample, we ensure that our estimates of the Instantaneous Carbon Feedback (ICF) and the structural lag effects are derived from pure physical signals. This rigorous selection process maximizes the internal validity of our causal identification strategy, as it eliminates the possibility that the identified null Instantaneous Greenhouse Effect (IGE) is a product of data smoothing or interpolation bias.
Figure 2: Visualization of the data gridding structure and observational gaps. The upper and middle panels show the density of 100-year and 1,000-year grid points for Temperature and CO2, respectively. The bottom panel identifies data gaps (missing observations). The period corresponding to Termination II is highlighted as the longest continuous interval free of observational gaps for both variables.
@@
Table 3: Comparison of VECM Results: Termination II vs. Whole Period
Termination II Whole Period
@@
Error Correction ($\alpha$) 0.103-0.103** $0.002$ 0.164-0.164*** 0.0002-0.0002
@@
Prior to estimating the structural dynamics of the climate system, we conduct a series of diagnostic tests to ensure the statistical validity of the Vector Error Correction framework. Table 1 reports the results of the Augmented Dickey-Fuller (ADF) unit root tests and the Engle-Granger cointegration tests. The ADF statistics indicate that both temperature and the logarithm of CO2 are non-stationary $I(1)$ processes across both the Whole Period and the Termination II samples. The Engle-Granger tests confirm the existence of a stable cointegrating relationship, providing the empirical basis for the VECM specification defined in Equation (1). Furthermore, the lag order selection criteria presented in Table 2 suggest that a second-order lag structure (where the level-VAR lag $k=3$ corresponds to VECM lag $p=2$) is optimal for capturing the millennial-scale dynamics while maintaining statistical parsimony.
The primary estimation results for the VECM are summarized in Table 3. Panel A presents the identified cointegrating vector ($\beta$), which defines the long-run physical bond between the variables. In both samples, we find a remarkably consistent and statistically significant equilibrium relationship. Specifically, for the Termination II sample—which provides the most reliable signal due to the absence of interpolation gaps—the relationship is estimated as $T-18.748\ln C=\textit{const}$. From this result, we derive the Earth System Sensitivity (ESS), defined as the total temperature response to a doubling of atmospheric CO2 concentration after all long-term feedbacks have reached equilibrium:
$\text{ESS}=18.748\times\ln(2)\approx 13.0\text{ K}$ (7)
Our estimate of 13.0 K is situated at the higher end of the range reported in the literature, such as the $9\pm 4$ K (95% CI: 7–13 K) estimate by Snyder (2016).
This high sensitivity can be understood through the conceptual framework of the PALAEOSENS project (Rohling et al., 2012). While the conventional Equilibrium Climate Sensitivity (ECS) typically accounts for “fast” feedbacks such as water vapor and clouds (often referred to as Charney sensitivity), our VECM, calibrated on the millennial-scale transitions of the Pleistocene, inherently captures the “slow” feedbacks of the Earth system. These include ice-sheet albedo changes, vegetation shifts, and long-term carbon-cycle responses that operate over multi-millennial time scales. According to Rohling et al. (2012), the inclusion of these slow feedbacks—which characterizes the transition from $S_{act}$ to $S_{ESS}$—can significantly amplify the total temperature response, potentially doubling the sensitivity compared to the Charney estimate. Our result of 13.0 K represents this comprehensive Earth System Sensitivity ($S_{ESS}$), reflecting the state-dependent amplification characteristic of major deglaciation events like Termination II.
Panel B of Table 3 delineates the short-run dynamics and the adjustment mechanisms of the system. The reduced-form lag coefficient for CO2 in the temperature equation, $(\Gamma_{1})_{12}=15.201$ ($p<0.01$), identifies the impact of CO2 changes from the previous millennium. Crucially, under the assumption of a null Instantaneous Greenhouse Effect ($a_{12}=0$)—a condition to be empirically validated in the subsequent section via the ICF Window for null IGE—this reduced-form parameter directly identifies the structural lag coefficient $(\Gamma_{1}^{*})_{12}$, representing the momentum of the greenhouse effect.
By recalling the structural identity $(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}_{1}\Delta y_{t-1}+\dots+u_{t}$, the relationship between the structural and reduced-form coefficients is defined by the identity $\Gamma^{*}_{1}=(I-A_{0})\Gamma_{1}$. The specific element for the CO2-to-temperature lag can be expanded as follows:
Owing to the strict exogeneity of orbital parameters, past climate states cannot influence current orbital forcing, which dictates that the reduced-form coefficient $(\Gamma_{1})_{32}$ is effectively zero. Under this physical constraint, and imposing the null-IGE identification ($a_{12}=0$), the second and third terms in the expansion vanish. Consequently, the structural lag effect is directly identified by the estimated reduced-form coefficient: $(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}=15.201$.
This structural identification allows for a precise decomposition of the total temperature response observed in the 1,000-year window following a CO2 doubling event ($\Delta\ln C=\ln 2\approx 0.693$). The resulting warming is the sum of two distinct physical forces:
  1. Structural Momentum: The structural lag effect $(\Gamma^{*}_{1})_{12}$ provides an immediate "kick" of $15.201\times\ln(2)\approx 10.5$ K. This represents the powerful, delayed radiative feedback that manifests after the initial 1,000-year period of relative stability defined by the null-IGE identification.
  2. Equilibrium Adjustment: The error correction coefficient $\alpha=-0.103$ ($p=0.013$) quantifies the system’s restoring force toward the long-run ESS. Starting from the disequilibrium created by the CO2 innovation, this mechanism contributes an additional adjustment of approximately $13.0\text{ K}\times 0.103\approx 1.3$ K.
Combined, these forces generate a total warming of approximately 11.8 K within the first millennium after the lag period. The fact that this combination of delayed structural momentum and equilibrium pull so effectively accounts for the millennial-scale temperature variance explains why the model identifies $a_{12}\approx 0$ at the contemporaneous scale. In our framework, the greenhouse effect is not diminished; rather, it is structurally localized as a delayed feedback, allowing the contemporaneous driver of climatic shifts to be identified as orbital-induced temperature changes.
4.1.1 Consistency with Modern Anthropogenic Warming
The robustness of our SVECM estimates is further validated by comparing the model’s projected warming rate with observed modern temperature trends. Our results for Termination II indicate that a doubling of CO2 generates a total millennial response of approximately 11.8 K, comprising the 10.5 K structural momentum and the 1.3 K equilibrium adjustment. This translates to a centennial warming rate of 1.18 K per 100 years for a CO2 doubling.
To contextualize this within the modern era, we consider the increase in atmospheric CO2 since the pre-industrial period, from a baseline of approximately 280 ppm to the current level of roughly 428 ppm. In logarithmic terms, this corresponds to an accumulation of approximately 61% of a doubling:
$\text{Modern Accumulation Factor}=\frac{\ln(428/280)}{\ln(2)}\approx 0.61$ (9)
Applying our derived centennial rate to this factor, the model predicts a CO2-induced warming of:
$1.18\text{ K/century}\times 0.61\approx 0.72\text{ K/century}$ (10)
Remarkably, this paleo-derived estimate is in near-perfect alignment with the observed instrumental warming rate of approximately 0.73–0.80 K per century (equivalent to the observed 1.1–1.2 K rise over the last 150 years).
This high degree of correspondence suggests that the null-IGE identification ($a_{12}=0$) at the millennial scale does not diminish the long-term radiative importance of CO2. Instead, it reveals that the intrinsic sensitivity of the Earth system—as captured during the rapid transitions of the Pleistocene—is sufficient to account for the bulk of modern warming through delayed structural feedbacks and equilibrium adjustments. The fact that a model calibrated on 130,000-year-old ice-core data can so accurately recover the scale of 21st-century climate change provides powerful empirical support for our causal identification strategy.
@@ 4.1.1 Consistency with Modern Anthropogenic Warming
4.2.1 Instantaneous Carbon Feedback (ICF)
The structural parameter $a_{21}$ is defined as a key element of the contemporaneous interaction matrix $A_{0}$ in our SVAR system, representing the simultaneous feedback within a single 1,000-year time step. Specifically, $a_{21}$ quantifies the Instantaneous Carbon Feedback (ICF), defined as the response of the logarithm of CO2 concentration to a change in temperature:
$a_{21}=\frac{\partial\ln C_{t}}{\partial T_{t}}$ (11)
@@ 4.2.1 Instantaneous Carbon Feedback (ICF)
$\Delta T_{t}=a_{12}\Delta\ln C_{t}+w_{t}$ (13)
where $a_{12}$ represents the Instantaneous Greenhouse Effect (IGE). If a significant IGE exists at the millennial resolution, the OLS estimate for $a_{21}$ would suffer from simultaneity bias, as $\Delta T_{t}$ would be correlated with the error term $v_{t}$.

### v3

# arXiv:2601.06986v3[physics.ao-ph] 28 Feb 2026

@@ Abstract
Using a gap-free, millennial-resolution record of atmospheric CO2 and Antarctic temperature since the Last Interglacial (130 ka ~), we structurally identify the feedback architecture of the climate–carbon system. The series are cointegrated, supporting estimation with a Vector Error Correction Model (VECM). The cointegrating relationship implies a temperature response of 13.0 K per CO2 doubling. Structural identification combining a Causality Allocation Curve (CAC) and Milankovitch-cycle instrumental variables yields a contemporaneous carbon response (CCR) of 5.77 ppm/K, while the contemporaneous temperature response (CTR) is statistically indistinguishable from zero. Accounting for lagged feedbacks, the cumulative temperature response within one millennium of CO2 doubling reaches 11.8 K.
@@ Introduction
The tight coupling between Antarctic temperature and atmospheric carbon dioxide (CO2) concentrations over the late Quaternary is one of the most striking features of the Earth system. While the high degree of correlation between these variables is well-established, the nature of their contemporaneous interaction—specifically the direction, magnitude, and timing of feedbacks at millennial resolutions—remains a subject of intense debate. Establishing a robust causal architecture is not only vital for understanding past glacial-interglacial transitions but also for constraining long-term Earth System Sensitivity (ESS) in the context of anthropogenic warming.
The statistical analysis of paleoclimate records faces three fundamental challenges. First, the high persistence of temperature and CO2 series carries the risk of spurious regression (Kaufmann et al., 2006; Cummins et al., 2022). Because these variables exhibit the properties of non-stationary, $I(1)$ processes, standard regression techniques can yield misleadingly significant correlations that lack physical causality. Second, a methodological schism exists regarding the appropriate framework for these series; while some advocate for a cointegration approach (Kaufmann et al., 2006; Miller, 2019), others argue that the system is ultimately mean-reverting and stationary, favoring standard Vector Autoregression (VAR) models (Davidson et al., 2016). Third, the irregular temporal spacing of ice-core proxies often necessitates linear interpolation, which has been shown to introduce substantial size distortions and biases in statistical testing (Miller, 2019).
Beyond these statistical hurdles, a long-standing climate-physical debate persists regarding the “lead-lag” relationship between temperature and CO2. Early studies suggested that CO2 lagged Antarctic temperature by several centuries during deglaciations (Caillon et al., 2003), implying that carbon was a follower rather than a driver of initial warming. More recent re-evaluations, accounting for gas-ice age uncertainties, have proposed a nearly synchronous relationship (Parrenin et al., 2013). However, these findings often rely on reduced-form correlations that cannot isolate structural innovations from reciprocal feedbacks. While Lemoine (2010) utilized orbital parameters as instrumental variables (IV) to identify a significantly positive—yet relatively modest—climate–carbon feedback factor, the short-run structural dynamics within a unified millennial-scale framework remain under-explored.
This study addresses these challenges by applying a Structural Vector Error Correction Model (SVECM) to a high-resolution, gap-free ice-core dataset. Our contribution is twofold. First, we identify a specific 130,000-year interval (the 130 ka sample) that is entirely free of observational gaps at a 1,000-year resolution. By utilizing this interval, we eliminate the need for interpolation, thereby physically neutralizing the statistical artifacts warned of by Miller (2019). Second, we employ a structural identification strategy that utilizes the strict exogeneity of Milankovitch cycles as an instrument to disentangle the contemporaneous temperature response (CTR, $a_{12}$) from the contemporaneous carbon response (CCR, $a_{21}$).
Our results reveal a stark asymmetry in the contemporaneous climate–carbon interaction. We find a robust, statistically significant CCR ($\hat{a}_{21}>0$), consistent with the magnitude suggested by last-millennium reconstructions (Frank et al., 2010). In contrast, the CTR ($a_{12}$) is found to be statistically indistinguishable from zero at the millennial scale. This identification implies that while temperature changes trigger an immediate carbon release, the radiative feedback of CO2 onto temperature requires structural lags to manifest. This framework provides a unified explanation for the perceived “synchronicity” in deglacial records as a manifestation of the CCR, while reconciling the “lag” observed in other contexts as the physical adjustment time of the thermal system. By bridging the gap between time-series econometrics and paleoclimatic physics, we provide a more precise characterization of the feedbacks that govern the Earth’s long-term climate sensitivity.
@@
To circumvent this risk while preserving the critical information contained in the levels of the variables, we employ a Vector Error Correction Model (VECM). This framework is appropriate when variables are cointegrated, meaning that while individual series may drift non-stationarily, a stable linear combination exists. In this study, we consider a three-variable system $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$, where $T_{t}$ is Antarctic temperature, $\ln C_{t}$ is the logarithm of CO2 concentration, and $M_{t}$ represents external orbital (Milankovitch) forcing. The VECM is expressed as:
@@
where $\beta$ contains the cointegrating vector, $\alpha$ represents the adjustment speeds to equilibrium, and $\Gamma_{i}$ captures the short-run lagged dynamics. The error term $\epsilon_{t}$ denotes the reduced-form innovations, which are generally correlated.
@@
To isolate structural shocks ($u_{t}$) from the reduced-form residuals ($\epsilon_{t}$), we define the relationship $(I-A_{0})\epsilon_{t}=u_{t}$, where $u_{t}$ is a vector of orthogonal structural innovations with $u_{t}\sim\text{i.i.d.}\mathcal{N}(0,\Sigma_{u})$. In our three-variable system, the contemporaneous interaction matrix $A_{0}$ is defined as:
$A_{0}=\begin{pmatrix}0&a_{12}&a_{13}\\ a_{21}&0&a_{23}\\ 0&0&0\end{pmatrix}$
where $a_{12}$ and $a_{21}$ represent the contemporaneous temperature response (CTR) and the contemporaneous carbon response (CCR), respectively. The third row is set to zero due to the strict exogeneity of orbital forcing, implying $\epsilon_{3t}=u_{3t}$. The structural equations for the climate residuals are then given by:
$u_{1t}$ $=\epsilon_{1t}-a_{12}\epsilon_{2t}-a_{13}\epsilon_{3t}$
$u_{2t}$ $=-a_{21}\epsilon_{1t}+\epsilon_{2t}-a_{23}\epsilon_{3t}$
By imposing the orthogonality condition $\mathbb{E}[u_{1t}u_{2t}]=0$ and noting that $\epsilon_{3t}$ is orthogonal to both $u_{1t}$ and $u_{2t}$ (since $\epsilon_{3t}=u_{3t}$), the terms involving orbital forcing effectively vanish in the expectation. This yields the Causality Allocation Curve (CAC), which defines the structural trade-off between the CTR and the CCR:
@@
Equation (2) demonstrates that the allocation of the total contemporaneous correlation between temperature and CO2 into directional causal links is uniquely determined by the sample covariance structure ($s_{ij}$) of the residuals.
@@
To visualize the propagation of these structural innovations, we utilize Impulse Response Functions (IRFs). The IRFs trace the time path of the system in response to a one-time shock in $u_{t}$. We simulate two primary scenarios:
@@
    A CO2 doubling shock: Represented as $u_{2t}=\ln(2)\approx 0.693$. This allows us to observe the transition from the transient response to the long-term Earth System Sensitivity (ESS) determined by $\beta$.
@@
    A 2 K temperature shock: Represented as $u_{1t}=2.0$. This isolates the magnitude and speed of the carbon cycle’s response via the CCR and the subsequent feedback loops.
The presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline. The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$ (see Appendix A, for details), provides a formal bridge between millennial-scale innovations and the long-term equilibrium states.
@@
Sources
@@
The empirical foundation of this study rests on three primary data series spanning the late Quaternary: atmospheric CO2 concentrations, Antarctic surface temperature anomalies, and solar insolation. Figure 1 illustrates the evolution of these variables over the past 800,000 years, highlighting the strong phase-locking of the climate–carbon system to orbital cycles.
Atmospheric CO2 data are obtained from the composite ice-core record (Lüthi et al., 2008; Bereiter et al., 2015), which integrates measurements from multiple Antarctic cores to capture natural greenhouse gas variability across eight complete glacial-interglacial cycles. For temperature estimates, we utilize the Antarctic temperature anomaly record ($\Delta T$) derived from deuterium ratios in the EPICA Dome C ice core (Jouzel et al., 2007), serving as a reliable indicator for global climatic shifts. External orbital forcing is represented by the solar insolation at 60N in June, computed based on the solution by Laskar et al. (2004).
Gridding and Sample Selection
To conduct a rigorous time-series analysis, the unevenly spaced raw data were transformed into a regular 1,000-year resolution grid. We distinguish between two primary sample periods: the comprehensive 800 ka sample and the focused 130 ka sample (covering the period from the Last Interglacial to the present).
Critically, as visualized in Figure 2, our grid analysis identifies that the 130 ka sample represents a continuous interval entirely free of observational gaps at the millennial resolution. Because this interval requires zero interpolation, the resulting series remains a direct reflection of the physical ice-core record, immune to statistical artifacts. By focusing on this gap-free 130 ka record, we ensure that the estimated structural parameters—particularly the CCR ($a_{21}$) and structural lag effects—reflect pure physical signals.
Figure 1: Evolution of climatic and orbital variables over the past 800,000 years. The panels display (from top to bottom) solar insolation at $60^{\circ}$N in June and December, Antarctic temperature anomalies ($\Delta T$), and the logarithm of atmospheric CO2 concentration ($\ln C$). The shaded area indicates the 130 ka sample period used for primary structural analysis.
Figure 2: Visualization of data gridding and observational gaps. The upper panels show the 100-year raw fragments, while the lower panels display the 1,000-year grid used as the VECM basis. The dashed box highlights the 130 ka sample as the longest continuous interval free of gaps for both temperature and CO2, justifying its use for structural identification.
@@
Table 3: Comparison of VECM Results: 130 ka vs. 800 ka samples
130 ka sample 800 ka sample
@@
Error Correction ($\alpha$) 0.103-0.103∗∗ $0.002$ 0.164-0.164∗∗∗ 0.0002-0.0002
@@
Prior to estimating the structural dynamics of the climate system, we conduct a series of diagnostic tests to ensure the statistical validity of the Vector Error Correction framework. Table 1 reports the results of the Augmented Dickey-Fuller (ADF) unit root tests and the Engle-Granger cointegration tests. The ADF statistics indicate that both temperature and the logarithm of CO2 are non-stationary $I(1)$ processes across both the 800 ka and the 130 ka samples. The Engle-Granger tests confirm the existence of a stable cointegrating relationship, providing the empirical basis for the VECM specification defined in Equation (1). Furthermore, the lag order selection criteria presented in Table 2 suggest that a second-order lag structure (where the level-VAR lag $k=3$ corresponds to VECM lag $p=2$) is optimal for capturing the millennial-scale dynamics while maintaining statistical parsimony.
The primary estimation results for the VECM, comparing the gap-free 130 ka sample with the interpolated 800 ka whole period, are summarized in Table 3. Panel A presents the estimated cointegrating vector ($\beta$), which defines the long-term equilibrium relationship between temperature and CO2. In both samples, the coefficient for $\ln C_{t-1}$ is negative and highly significant ($p<0.01$), indicating a robust positive long-term correlation between climatic warming and atmospheric carbon concentration.
Panel B describes the short-run dynamics and the error correction mechanism. The error correction term ($\alpha$) for the temperature equation is negative and statistically significant in both samples (-0.103 for the 130 ka sample), confirming that temperature deviations from the long-term carbon-linked equilibrium are corrected over time. In contrast, the error correction term for the $\ln C_{t}$ equation is statistically indistinguishable from zero, suggesting that atmospheric CO2 acts as the primary stochastic driver of the long-term trend in this system. The short-run lagged effects ($\Delta T_{t-i},\Delta\ln C_{t-i}$) also reveal significant millennial-scale persistence and cross-variable interactions. These estimated parameters provide the necessary foundation for the structural causality analysis and sensitivity evaluations performed in the subsequent sections.
@@
The structural parameter $a_{21}$ is defined as a key element of the contemporaneous interaction matrix $A_{0}$ in our SVAR system, representing the simultaneous feedback within a single 1,000-year time step. Specifically, $a_{21}$ quantifies the contemporaneous carbon response (CCR), defined as the response of the logarithm of CO2 concentration to a change in temperature:
$a_{21}=\frac{\partial\ln C_{t}}{\partial T_{t}}$
@@
$\Delta T_{t}=a_{12}\Delta\ln C_{t}+w_{t}$
where $a_{12}$ represents the contemporaneous temperature response (CTR). If a significant CTR exists at the millennial resolution, the OLS estimate for $a_{21}$ would suffer from simultaneity bias, as $\Delta T_{t}$ would be correlated with the error term $v_{t}$.